# BAD-NEUS: Rapidly converging trajectory stratification

### v1

@@ Abstract
An issue for molecular dynamics simulations is that events of interest often involve timescales that are much longer than the simulation time step, which is set by the fastest timescales of the model. Because of this timescale separation, direct simulation of many events is prohibitively computationally costly. This issue can be overcome by aggregating information from many relatively short simulations that sample segments of trajectories involving events of interest. This is the strategy of Markov state models (MSMs) and related approaches, but such methods suffer from approximation error because the variables defining the states generally do not capture the dynamics fully. By contrast, once converged, the weighted ensemble (WE) method aggregates information from trajectory segments so as to yield unbiased estimates of both thermodynamic and kinetic statistics. Unfortunately, errors decay no faster than unbiased simulation in WE. Here we introduce a theoretical framework for describing WE that shows that introduction of an element of stratification, as in nonequilibrium umbrella sampling (NEUS), accelerates convergence. Then, building on ideas from MSMs and related methods, we propose an improved stratification that allows approximation error to be reduced systematically. We show that the improved stratification can decrease simulation times required to achieve a desired precision by orders of magnitude.
@@ I Introduction
The sampling of a stochastic process can be controlled by evolving an ensemble and splitting and merging the trajectories of its members, and various algorithms based on this strategy have been introduced [1, 2, 3, 4, 5, 6, 7]. Because the trajectory segments between splitting events are unbiased, such algorithms can yield dynamical statistics, such as transition probabilities and mean first passage times (MFPTs) to selected states, in addition to steady-state probability distributions and averages. Furthermore, splitting algorithms generally require little communication between the members of the ensemble, making them relatively straightforward to implement regardless of the underlying dynamics, and community software is available [8]. As a result, splitting algorithms are widely used.
@@
Others focused on improving the rate estimates from WE [12, 13]. Of particular relevance to the present study, Copperman and Zuckerman exploited the fact that splitting algorithms sample unbiased trajectory segments to construct history-augmented Markov state models (haMSMs) [14] from their WE data [12]. In an haMSM, the system is divided into discrete states and the probabilities for transitions between them are computed from data, just as in a standard MSM, except that the path ensemble is split based on the last metastable state visited; the rate constant is then computed as the flux from $A$ to $B$ conditioned on having last visited $A$, normalized by the steady-state probability of having last visited $A$. The idea is that finer discretization of the state space than the WE simulation yields improved estimates. However, this approach does not accelerate the convergence of the sampling.
@@
The basic approach in our work is to combine these strategies and fully integrate them in the sampling algorithm by applying a short-trajectory based approximation to the stationary distribution between successive iterations. As we show below, if the approximation satisfies certain properties and is correctly integrated in the algorithm, we can accelerate the convergence by several orders of magnitude without introducing systematic bias. Our approach is able to compute the full stationary distribution with minimal assumptions about the dynamics, as well as kinetic statistics such as rates and committors (probabilities of visiting selected states before others).
@@
Conceptually, one can view our approach as a form of stratification, in which the weights of not just individual trajectory segments but groups of them are manipulated. Stratification was first introduced for controlling the sampling of trajectory segments in nonequilibrium umbrella sampling (NEUS) [3, 15, 16, 4, 17, 18]. NEUS groups trajectories that are in the same regions of state space, as defined by collective variables (CVs), and estimates the steady-state probabilities of the regions by solving a global flux balance equation; this strategy was subsequently adopted in extensions of WE[19] and exact milestoning (EM)[6]. A unified framework for trajectory stratification that incorporates elements of all of the above algorithms and is unbiased in the limit that each region contains an infinite number of ensemble members is presented in Ref. 18, which also shows that the regions can be defined in terms of path-based quantities. Mathematical analysis of trajectory stratification algorithms can be found in Refs. 20, 21.
@@
Our paper is organized as follows. In Section II, we review the WE algorithm and recast it to show why an initialization bias is slow to disappear. This formulation also clarifies the relation of WE to the NEUS algorithm as described in Ref. 18, and we show how software for WE can be easily extended to enable trajectory stratification in Section III. In Section IV, we introduce a generalization of MSMs that represents the dynamics through a basis expansion [22, 23, 24] and in turn our strategy for accelerating sampling, which we term Basis-Accelerated NEUS (BAD-NEUS). We show that the NEUS algorithm is a special case of our new scheme. In Sections V and VI, we demonstrate BAD-NEUS on a two-dimensional model with a known steady-state distribution and a molecular example.
@@ II Weighted Ensemble (WE)
As described above, in WE, the state space is partitioned into regions, and then trajectories are copied (cloned or split) or removed (killed or pruned) from the ensemble based on criteria. Mathematically, we denote the state of ensemble member (henceforth, walker) $i$ at time $t$ by $X^{i}_{t}$, and we associate with $i$ a weight $w^{i}_{t}$ such that $\sum_{i}w^{i}_{0}=1$. We track the index of the region containing $X^{i}_{t}$ by an index process $J^{i}_{t}$. For example, $J^{i}_{t}$ might track the element of a partition of state space in which $X^{i}_{t}$ currently resides. Other options for the index process are available and can be advantageous (see Refs. 18, 25 for further discussion).
@@
where the subscript on the expectation indicates that we draw the initial state $X_{0}$ from the steady-state distribution $\pi,$ and $n$ indexes successive weighted ensemble iterations. We present necessary modifications for the case where $\Delta$ is a random variable later. Here and below, $X$ without a superscript indicates a realization of the underlying Markov process rather than a walker in an ensemble. An important example is the steady-state flux into a set $D$ of interest, which can be used to compute the MFPT to $D$ from another set in the right algorithmic setup [9]. This flux can be obtained by setting $g(X_{0},\ldots,X_{\Delta})=\mathbbm{1}_{D^{c}}(X_{0})\mathbbm{1}_{D}(X_{1})$, where $\mathbbm{1}_{D}(x)$ is an indicator function that is 1 if $x\in D$ and 0 otherwise. This amounts to counting the number of walkers that enter $D$ in a single time step and summing the total weight of those walkers.
@@
where $dx$ is an infinitesimal volume in state space that is located at a specific value of $x$ and $S(\ell)$ is an increasing sequence of stopping times. For example, if we choose $S(\ell)=\ell\Delta$ for a fixed $\Delta$, the following description corresponds to the standard version of WE already sketched. Alternatively, we can let $S(\ell)$ be the time of the $\ell$-th change in the value of the index process $J$:
@@
The outputs of the weighted ensemble iteration and the accelerated variants that we consider here are approximations of the steady-state conditional flux distributions, as well as estimates of the corresponding normalization constants. The conditional flux distributions are related to joint flux distributions through the normalization constants:
@@
The subscript on the weight does not change because the evolution step only affects the state and index of a walker, not its weight. $\tilde{\mathcal{U}}$ as defined in (10) is an unbiased stochastic approximation of $\mathcal{U}$ in the sense that, for any function $g$ and any distribution $\rho$,
@@
If distributions are represented with some ansatz such as a neural network, one can also apply an approximate propagator based on a variational method as outlined in Refs. 26, 27. Such a scheme would not require new trajectories to be run at each iteration, nor would it require a resampling step.
As noted above, we also represent resampling through an operator, $\mathcal{R}_{k}$. We define it such that, for any distribution $\rho$ and any function $g$
@@
Above, the LHS selects the portion of the distribution $\rho$ in region $k$ and then renormalizes it, while the RHS corresponds to resampling from the distribution. A simple choice which is commonly employed for the operator $\mathcal{R}_{k}$ is to sample the set of walkers in region $k$ at the end of the evolution (denoted $K$) from a multinomial distribution with probability proportional to $\{w_{S^{i}(\ell)}^{i}\mathbbm{1}_{\{k\}}(J_{S^{i}(\ell)}^{i})\}_{i\in K}$ with $N_{k}$ trials and then return the distribution
@@
In words, (14) shows that ${\cal R}_{k}$ constructs the state distribution from a sum over the resampled walkers in each region. There are other possible ways of resampling in the walker representation, such as stratified or pivotal resampling [28], and these may in practice be better. However, any choice that satisfies (12) suffices for our discussion.
@@
Algorithm 1 suggests that, if $\tilde{\overline{\pi}}=\overline{\pi}$ in step 3, convergence would be achieved. Therefore, our strategy for accelerating WE is to replace the power iteration step with a more accurate approximation. Specifically, we compute changes of measure $\nu_{\ell}$ such that
@@ III Nonequilibrium umbrellla sampling (NEUS)
We now present NEUS as a simple extension to WE that accelerates convergence in the way suggested in (17); EM [6] can be formulated similarly. Our development is based on the algorithm in Ref. 18, which corrects a small systematic bias in earlier NEUS papers [3, 15, 16, 4, 17]. We begin by making the observation that, at steady state,
@@
$\int\overline{\pi}(dx,k)=\sum_{j}\int\overline{\pi}(dx,j)\mathbbm{E}_{X_{0}=x,J_{0}=j}[\mathbbm{1}_{\{k\}}(J_{S(1)})].$ (19)
This observation introduces one equation per possible value of the index $j$. Therefore, we can parameterize the change of measure with as many free parameters:
@@
$\overline{G}_{jk}=\int\overline{\pi}(dx|j)\mathbbm{E}_{X_{0}=x,J_{0}=j}[\mathbbm{1}_{\{k\}}(J_{S(1)})].$ (22)
We summarize NEUS in Algorithm 2.
@@
We thus see that NEUS is distinguished from WE by steps 2 and 3. In practice, walkers are drawn and their dynamics are simulated to determine $S(\ell+1)$ in the loop over $\ell$ prior to step 2 (the resulting state distribution is used later in step 5); then, $\overline{G}_{jk}$ is computed in step 2 from walkers at iteration $\ell$ as
@@
We note that, if the flux distributions $\overline{\pi}(dx|k)$ and normalization constants $\overline{z}^{k}$ assume their steady-state values, then (21) is satisfied by $c^{k}=1$, and NEUS has the correct fixed point.
@@ IV BAD-NEUS
We then introduce the basis set $\{\phi_{p}(x,k)\}_{p}$, and write
@@
$0=\sum_{kp}c_{\ell}^{p}\overline{z}_{\ell}^{k}\int\phi_{p}(x,k)\overline{\pi}_{\ell}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}\phi_{r}(X_{t},J_{t})-\phi_{r}(X_{t+\tau},J_{t+\tau})\right].$ (27)
@@
The basis set must include the constant function in its span so that the system has a unique nontrivial solution. The NEUS algorithm corresponds to the choice
$\phi_{p}(x,k)=\mathbbm{1}_{\{p\}}(k)$ (30)
and $\tau=1$. We summarize the BAD-NEUS algorithm in Algorithm 3.
$M_{pr}\approx\sum_{ki}\overline{z}_{\ell}^{k}\phi_{p}(X^{i}_{S^{i}(\ell)},k)\mathbbm{1}_{\{k\}}(J^{i}_{0})\left[\sum_{t=S^{i}(\ell)}^{S^{i}(\ell+1)-1}(\phi_{r}(X^{i}_{t},J_{t}^{i})-\phi_{r}(X^{i}_{t+\tau},J_{t+\tau}^{i}))\right].$ (31)
where $Z$ normalizes the total of the walker weights to one. If the basis set contains the constant function in its span, and the flux distributions $\overline{\pi}_{\ell}(dx|k)$ and region weights $\overline{z}_{\ell}^{k}$ take their steady-state values, then the linear system (27) is solved by $\sum_{p}c_{p}\phi(x,k)=1$, and BAD-NEUS has the correct fixed point. To reduce variance in our estimate of the matrix $M$ to ensure a stable solve, it is often useful to work with the mixture distribution for the last $h$ iterations:
Finally, we note that the general strategy of improving the approximation of the change of measure is not limited to using a basis expansion. Just as one can use various means to solve the equations of the operator that encodes the statistics of the dynamics [22, 23, 29, 27], one can represent the change of measure here by either a basis expansion or a nonlinear representation. In particular, our tests based on the neural-network approach in Ref. 27 show significant promise (unpublished results).
@@
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, Resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$, basis set $\{\phi_{p}(x,k)\}_{p}$, lag time $\tau$

### v2

@@ Abstract
An issue for molecular dynamics simulations is that events of interest often involve timescales that are much longer than the simulation time step, which is set by the fastest timescales of the model. Because of this timescale separation, direct simulation of many events is prohibitively computationally costly. This issue can be overcome by aggregating information from many relatively short simulations that sample segments of trajectories involving events of interest. This is the strategy of Markov state models (MSMs) and related approaches, but such methods suffer from approximation error because the variables defining the states generally do not capture the dynamics fully. By contrast, once converged, the weighted ensemble (WE) method aggregates information from trajectory segments so as to yield unbiased estimates of both thermodynamic and kinetic statistics. Unfortunately, errors decay no faster than unbiased simulation in WE. Here we introduce a theoretical framework for describing WE that shows that introduction of an approximate stationary distribution on top of the stratification, as in nonequilibrium umbrella sampling (NEUS), accelerates convergence. Building on ideas from MSMs and related methods, we generalize the NEUS approach in such a way that the approximation error can be reduced systematically. We show that the improved algorithm can decrease simulation times required to achieve a desired precision by orders of magnitude.
@@ I Introduction
The sampling of a stochastic process can be controlled by evolving an ensemble and splitting and merging the trajectories of its members, and various algorithms based on this strategy have been introduced [1, 2, 3, 4, 5, 6, 7]. Because the trajectory segments between splitting events are unbiased, such algorithms can yield dynamical statistics, such as transition probabilities and mean first passage times (MFPTs) to selected states, in addition to steady-state probability distributions and averages. Furthermore, splitting algorithms generally require little communication between the members of the ensemble, making them relatively straightforward to implement regardless of the underlying dynamics, and community software is available [8, 9]. As a result, splitting algorithms are widely used.
@@
Others focused on improving the rate estimates from WE [13, 14]. Of particular relevance to the present study, Copperman and Zuckerman exploited the fact that splitting algorithms sample unbiased trajectory segments to construct history-augmented Markov state models (haMSMs) [15] from their WE data [13]. In an haMSM, the system is divided into discrete states and the probabilities for transitions between them are computed from data, just as in a standard MSM, except that the path ensemble is split based on the last metastable state visited; the rate constant is then computed as the flux from $A$ to $B$ conditioned on having last visited $A$, normalized by the steady-state probability of having last visited $A$. The idea is that finer discretization of the state space than the WE simulation yields improved estimates. In itself, this approach does not accelerate the convergence of the sampling. However, Refs. 13, 8 suggest that iteratively restarting the WE simulation from the steady-state distribution estimated from the haMSM could accelerate convergence.
@@
Here, we build on this idea to introduce a general framework for how to apply a short-trajectory based approximation to the stationary distribution between successive iterations. As we show below, the approximation must satisfy certain properties and be correctly integrated in the algorithm; if these criteria are met, the approach can accelerate the convergence by several orders of magnitude without introducing systematic bias. Our approach is able to compute the full stationary distribution with minimal assumptions about the dynamics, as well as kinetic statistics such as rates and committors (probabilities of visiting selected states before others).
@@
Conceptually, one can view our approach as a form of stratification, in which the weights of not just individual trajectory segments but groups of them are manipulated. Stratification was first introduced for controlling the sampling of trajectory segments in nonequilibrium umbrella sampling (NEUS) [3, 16, 17, 4, 18, 19]. NEUS groups trajectories that are in the same regions of state space, as defined by collective variables (CVs), and estimates the steady-state probabilities of the regions by solving a global flux balance equation; this strategy was subsequently adopted in extensions of WE [20] and exact milestoning (EM)[6]. A unified framework for trajectory stratification that incorporates elements of all of the above algorithms and is unbiased in the limit that each region contains an infinite number of ensemble members is presented in Ref. 19, which also shows that the regions can be defined in terms of path-based quantities. Mathematical analysis of trajectory stratification algorithms can be found in Refs. 21, 22.
@@
Our paper is organized as follows. In Section II, we review the WE algorithm and recast it to show why an initialization bias is slow to disappear. This formulation also clarifies the relation of WE to the NEUS algorithm as described in Ref. 19, and we show how software for WE can be easily extended to enable trajectory stratification in Section III. In Section IV, we introduce a generalization of MSMs that represents the dynamics through a basis expansion [23, 24, 25] and in turn our strategy for accelerating sampling, which we term Basis-Accelerated NEUS (BAD-NEUS). We show that the NEUS algorithm is a special case of our new scheme. In Sections V and VI, we demonstrate BAD-NEUS on a two-dimensional model with a known steady-state distribution and a molecular example.
@@ II Weighted Ensemble (WE)
As described above, in WE, the state space is partitioned, and then trajectories are copied (cloned or split) or removed (killed or pruned) from the ensemble based on criteria. Mathematically, we denote the state of ensemble member (henceforth, walker) $i$ at time $t$ by $X^{i}_{t}$, and we associate with $i$ a weight $w^{i}_{t}$ such that $\sum_{i}w^{i}_{0}=1$. We track the index of the region containing $X^{i}_{t}$ by an index process $J^{i}_{t}$. For example, $J^{i}_{t}$ might track the element of a partition of state space in which $X^{i}_{t}$ currently resides. However, much more general choices, such as ones based on path quantities [19, 26] or distances between walkers [27, 28], are also possible. See Ref. 19 for further discussion.
@@
where the subscript on the expectation indicates that we draw the initial state $X_{0}$ from the steady-state distribution $\pi,$ and $n$ indexes successive weighted ensemble iterations. We present necessary modifications for the case where $\Delta$ is a random variable later. Here and below, $X$ without a superscript indicates a realization of the underlying Markov process rather than a walker in an ensemble. An important example is the steady-state flux into a set $D$ of interest, which can be used to compute the MFPT to $D$ from another set in the right algorithmic setup [10]. This flux can be obtained by setting $g(X_{0},\ldots,X_{\Delta})=\mathbbm{1}_{D^{c}}(X_{0})\mathbbm{1}_{D}(X_{1})$, where $\mathbbm{1}_{D}(x)$ is an indicator function that is 1 if $x\in D$ and 0 otherwise, and $D^{c}$ is the complement of the set $D$. This amounts to counting the number of walkers that enter $D$ in a single time step and summing the total weight of those walkers.
@@
where $dx$ is an infinitesimal volume in state space that is located at a specific value of $x$, and $S(\ell)$ is an increasing sequence of stopping times. For example, if we choose $S(\ell)=\ell\Delta$ for a fixed $\Delta$, the following description corresponds to the standard version of WE already sketched. Alternatively, we can let $S(\ell)$ be the time of the $\ell$-th change in the value of the index process $J$:
@@
This choice of stopping time has the advantage that walkers cannot go significantly beyond their regions, providing more control over the sampling.
@@
The outputs of the weighted ensemble iteration and the accelerated variants that we consider here are approximations of the steady-state conditional flux distributions, as well as estimates of the corresponding normalization constants (region weights). The conditional flux distributions are related to joint flux distributions through the normalization constants:
@@
The subscript on the weight does not change because the evolution step only affects the state and index of a walker, not its weight. Operationally, this equation corresponds to propagating an ensemble’s empirical distribution by running the dynamics until the stopping criterion, and then assigning the initial weight of each walker to its final state and index. $\tilde{\mathcal{U}}$ as defined in (10) is an unbiased stochastic approximation of $\mathcal{U}$ in the sense that, for any function $g$ and any distribution $\rho$,
@@
If distributions are represented with some ansatz such as a neural network, one can also apply an approximate propagator based on a variational method as outlined in Refs. 29, 30. Such a scheme would not require new trajectories to be run at each iteration, nor would it require a resampling step.
As noted above, we represent resampling through an operator, $\mathcal{R}_{k}$. We define it such that, for any distribution $\rho$ and any function $g$
@@
Above, the LHS of (12) selects the portion of the distribution $\rho$ in region $k$ and then renormalizes it, while the RHS corresponds to resampling from the distribution. This equation expresses the condition that the resampling operator must preserve the weighted distribution in the sense that the empirical mean of any function is the same in expectation after resampling. A simple choice which is commonly employed for the operator $\mathcal{R}_{k}$ is to sample the set of walkers in region $k$ at the end of the evolution (denoted $K$) from a multinomial distribution with probability proportional to $\{w_{S^{i}(\ell)}^{i}\mathbbm{1}_{\{k\}}(J_{S^{i}(\ell)}^{i})\}_{i\in K}$ with $N_{k}$ trials and then return the distribution
@@
In words, (14) shows that ${\cal R}_{k}$ constructs the state distribution from a sum over the resampled walkers in each region. There are other possible ways of resampling in the walker representation [31, 27, 28], and these may in practice be better. However, any choice that satisfies (12) suffices for our discussion.
@@
Algorithm 1 suggests that, if $\tilde{\overline{\pi}}=\overline{\pi}$ in step 3, convergence would be achieved. Therefore, our strategy for accelerating WE is to replace the power iteration step with a more accurate approximation. Specifically, we compute changes of measure (i.e., reweighting factors) $\nu_{\ell}$ such that
@@ III Nonequilibrium umbrellla sampling (NEUS)
We now present NEUS as a simple extension to WE that accelerates convergence in the way suggested in (17); EM [6] can be formulated similarly. Our development is based on the algorithm in Ref. 19, which corrects a small systematic bias in earlier NEUS papers [3, 16, 17, 4, 18]. We begin by making the observation that, at steady state,
@@
$\int\overline{\pi}(dx,k)=\\ \sum_{j}\int\overline{\pi}(dx,j)\mathbbm{E}_{X_{0}=x,J_{0}=j}[\mathbbm{1}_{\{k\}}(J_{S(1)})].$ (19)
This equation says that the steady-state probability of index process $k$ is the same as the total probability of walkers that were initialized from the steady-state distribution and flowed into (or remained in) $k$ at the stopping time. This observation introduces one equation per possible value of the index $k$. Therefore, we can parameterize the change of measure with as many free parameters:
@@
$\overline{G}_{jk}=\int\overline{\pi}(dx|j)\mathbbm{E}_{X_{0}=x,J_{0}=j}[\mathbbm{1}_{\{k\}}(J_{S(1)})]$ (22)
tracks the total flux from region $j$ to region $k$. We summarize NEUS in Algorithm 2.
@@
We thus see that NEUS is distinguished from WE in its original formulation by steps 2 to 4, in which the total fluxes between pairs of regions are estimated, and the region weights are adjusted to satisfy a global balance condition. In practice, walkers are drawn and their dynamics are simulated to determine $S(\ell+1)$ in the loop over $\ell$ prior to step 2 (the resulting state distribution is used later in step 5); then, $\overline{G}_{jk}$ is computed in step 2 from walkers at iteration $\ell$ as
@@
If the flux distributions and region weights assume their steady-state values such that $\overline{\pi}_{\ell}(dx|k)=\overline{\pi}(dx|k)$ and $\overline{z}_{\ell}^{k}=\overline{z}^{k}$, a global balance condition is satisfied:
$\overline{z}^{k}=\sum_{j}\overline{z}^{j}\overline{G}_{jk}.$ (24)
In this limit, (21) is solved by $c^{k}=1$, and NEUS has the correct fixed point (i.e., the same fixed point as $\cal U$, which encodes the unbiased dynamics).
@@ IV BAD-NEUS
We then introduce the basis set $\{\phi_{p}(x,k)\}_{p}$ for the change of measure in (17), and write
@@
$0=\sum_{k}\sum_{p}c_{\ell}^{p}\overline{z}_{\ell}^{k}\int\phi_{p}(x,k)\overline{\pi}_{\ell}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}\phi_{r}(X_{t},J_{t})-\phi_{r}(X_{t+\tau},J_{t+\tau})\right].$ (28)
@@
The basis set must include the constant function in its span so that the system has a unique nontrivial solution.
@@
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, Resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$, basis set $\{\phi_{p}(x,k)\}_{p}$, lag time $\tau$, total number of iterations $L.$
@@
We summarize the BAD-NEUS algorithm in Algorithm 3. If the flux distributions and region weights assume their steady-state values such that $\overline{\pi}_{\ell}(dx|k)=\overline{\pi}(dx|k)$ and $\overline{z}_{\ell}^{k}=\overline{z}^{k}$, then (26) is satisfied for all $f$. In that case, just as (24) and (21) lead to $c^{k}=1$ in NEUS, (26) and (28) lead to $\sum_{p}c_{p}\phi(x,k)=1$ in BAD-NEUS. In this case, step 6 in Algorithm 3 reduces to $\overline{\pi}_{\ell+1}(dx,k)\leftarrow\overline{z}^{k}\overline{\pi}(dx|k)$, and BAD-NEUS has the correct fixed point.
Specific choices for the basis functions, lag time, and stopping time correspond to existing methods. NEUS as described in Section III and Algorithm 2 corresponds to the choice
$\phi_{p}(x,k)=\mathbbm{1}_{\{p\}}(k),$ (31)
$\tau=1$, and $\Delta$ as defined in (3). We can obtain a version of iteratively restarted WE as suggested in Refs. 13, 8 by choosing the basis set to be indicator functions that are one on sets that are finer than the stratification (“microbins”), $\tau=1$, and a fixed stopping time $\Delta$. For these choices, the sum over $t$ in (30) telescopes, and $M$ can be row-normalized and written as $T-I$, where $T$ is the transition matrix for the MSM defined by the basis set.
$M_{pr}\approx\sum_{k}\sum_{i}\overline{z}_{\ell}^{k}\phi_{p}(X^{i}_{S^{i}(\ell)},k)\mathbbm{1}_{\{k\}}(J^{i}_{0})\left[\sum_{t=S^{i}(\ell)}^{S^{i}(\ell+1)-1}(\phi_{r}(X^{i}_{t},J_{t}^{i})-\phi_{r}(X^{i}_{t+\tau},J_{t+\tau}^{i}))\right].$ (32)
where $Z$ normalizes the total of the walker weights to one. Operationally, we loop over each walker before the resampling step and evaluate (33) to adjust its weight. Physically, this corresponds to apportioning the weight of the region containing the walker ($\overline{z}_{\ell}^{J_{S^{i}(\ell)}^{i}}/Z$ in (33)) according to the basis expansion representing the change of measure to the steady-state. We then proceed with the resampling step as in the usual WE algorithm.
@@
To reduce variance in our estimate of the matrix $M$ to ensure a stable solve, it is often useful to work with the mixture distribution for the last $h$ iterations:
Finally, we note that the general strategy of improving the approximation of the change of measure is not limited to using a basis expansion. Just as one can use various means to solve the equations of the operator that encodes the statistics of the dynamics [23, 24, 32, 30], one can represent the change of measure here by either a basis expansion or a nonlinear representation. In particular, our tests based on the neural-network approach in Ref. 30 show significant promise (unpublished results).